A jar of tea is placed in sunlight until it reaches an equilibrium temperature of 30.7 ◦C . In an attempt to cool the liquid, which has a mass of 187 g , 130 g of ice at 0.0 ◦C is added. At the time at which the temperature of the tea is 24.6 ◦C , find the mass of the remaining ice in the jar. The specific heat of water is 4186 J/kg · ◦ C . Assume the specific heat capacity of the tea to be that of pure liquid water. Answer in units of g. 


1
Expert's answer
2021-05-10T16:35:22-0400

Solution.

"m_t=187g=0.187kg;"

"t_t=30.7^oC;"

"m_i=130g=0.130kg;"

"t_i=0^oC;"

"t=24.6^oC;"

"c=4186J\/kg^oC;"

"\lambda=332000J/kg;"

"c_tm_t(t_t-t)=\lambda m+c_wm(t-t_i);\implies m=\dfrac{c_tm_t(t_t-t)}{\lambda +c_w(t-t_i)};"

"m=\dfrac{4186J/kg^oC\sdot0.187kg\sdot(30.7^oC-24.6^oC)}{332000J/kg+4186J/kg^oC\sdot(24.6^oC-0^oC)}=0.011kg=11g;"

Answer: "m=11g."